OpenStudy (anonymous):

21x^2+46x+24 I think the answer is (7x+6)(3x+4) If I am not correct why?

OpenStudy (anonymous):

For the x^2 part it should factor into 7x * 3x or 21x * x Let's try 7x * 3x: (7x ) (3x ) For the 24 part there are a bunch of possibilities: 1 * 24, 2 * 12, 3 * 8, 4 * 6 Let's try 3 * 8 (7x 3) (3x 8) 7*8 = 56, 3*3 = 9 Can't make 46 out of 56 and 9. . . (7x 8) (3x 3) 7*3 = 21, 8*3 = 24 Can't make 46 out of 21 and 24. . . Let's try 4 and 6 (7x 4) (3x 6) 7*6 = 42, 4*3 = 12 Can't make 46 out of 42 and 12 (7x 6) (3x 4) 7*4 = 28, 6*3 = 18 We can make 46 out of 28 and 18, so let's try this: (7x + 6) (3x + 4) Expand that out and we get. . . your right